OpenStudy (anonymous):

What is the simplest form of the number? 16^3/4

OpenStudy (shamim):

tell me 16^3=?

OpenStudy (paki):

16*16*16 / 4 ... simplify now...

OpenStudy (anonymous):

16*16*16=4096 4096/4=1024

OpenStudy (anonymous):

@shamim @paki

OpenStudy (paki):

yeah 1024... good...

OpenStudy (anonymous):

the equation looks like this \[16^{3/4}\]

OpenStudy (paki):

hmm

OpenStudy (paki):

then take (2)^4*3/4.... now cancel the 4 and you will be left with (2)^3... and take 2*2*2...

OpenStudy (paki):

clear now...?

OpenStudy (paki):

@ChasingMaggie

OpenStudy (anonymous):

so it's just 8? @paki

OpenStudy (paki):

yeah...

OpenStudy (anonymous):

thanks! Cna you help me with other problems?

OpenStudy (paki):

yeah i will try... and close this question now please...

OpenStudy (anonymous):

What is the solution to \[\sqrt{2x-1}=3\]

OpenStudy (paki):

take square on both sides... it will cancel the root... and then solve for "x"...

OpenStudy (anonymous):

So it's just \[2x-1=3\]

OpenStudy (paki):

no... 2x-1=9... now solve for "x"...

OpenStudy (anonymous):

x=5?

OpenStudy (paki):

yup

OpenStudy (paki):

close this question now... please...

OpenStudy (anonymous):

What is the solution to \[x=\sqrt{-x+6}\]

OpenStudy (anonymous):

@paki

OpenStudy (paki):

same as the previous one....

OpenStudy (anonymous):

well how can you square x? That's what's tripping me up

OpenStudy (anonymous):

Is it -3? @paki

OpenStudy (paki):

what is your question...?